# Sensing with near-infrared laser trapped fluorescent nanodiamonds

### v1

@@ Abstract
Biosensing based on optically trapped fluorescent nanodiamonds is an intriguing research direction potentially allowing to resolve biochemical processes inside living cells. Towards this goal, we investigate infrared near (NIR) laser irradiation at 1064 nm on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers. By conducting comprehensive experiments, we aim to understand how NIR exposure influences the fluorescence and sensing properties of FNDs and to determine the potential implications for the use of FNDs in various sensing applications. The experimental setup involved exposing FNDs to varying intensities of NIR laser light and analyzing the resultant changes in their optical and physical properties. Key measurements included T1 relaxation times, optical spectroscopy, and optically detected magnetic resonance (ODMR) spectra. The findings reveal how increased NIR laser power correlates with alterations in ODMR central frequency but also that charge state dynamics under NIR irradiation of NV centers plays a role. We suggest protocols with NIR and green light that mitigate the effect of NIR, and demonstrate that FND biosensing works well with such a protocol.
@@
Figure 1: Graphical abstract illustrating the exploration of effects of NIR laser on NV center for bio-sensing.
@@ I Introduction
Fluorescent nanodiamonds (FNDs) containing Nitrogen-Vacancy (NV) centers have attracted significant interest due to their combined optical and magnetic properties, which render them highly versatile for applications in particular in biomedical sensing and imaging [1, 2]. The capacity of NV centers for optical polarization and readout of spin states earmarks them as formidable tools for advanced sensing applications [3]. Yet, the utility of these sensors is heavily influenced by the photodynamics of the NV centers’ spin- and charge-states, which in turn governs their sensitivity and accuracy [4].
NV centers possess an electronic spin S=1 system, characterized by spin sublevels $m_{s}=0$ and $m_{s}=\pm 1$, which are separated by a zero-field splitting due to spin-spin interaction. This zero-field splitting is sensitive to external parameters such as temperature, strain, and electric fields, while the energies of the spin sublevels can be further shifted by magnetic fields through the Zeeman effect [5, 6]. Additionally, the optical spectrum of NV centers, including their fluorescence intensity and wavelength, is determined by the charge state distribution and depends on the local electrochemical potential and intrinsic charge state dynamics. These dependencies allow NV centers to act as versatile nanoscale sensors, providing insights into local magnetic fields across various frequency ranges, temperature fluctuations, and surface electrochemical conditions. Therefore, to accurately measure external parameters like magnetic fields or temperature with high sensitivity in biological environments, it is essential to understand how external control—such as a near infrared trapping laser—affect the intrinsic properties of NV centers [7, 8].
Trapping of FNDs with optical tweezers is an attractive approach as it potentially allows for spatially resolved sensing in challenging environments, with a resolution determined by the size of the FND and of the trapping potential [9]. To achieve this in biological samples and minimize cell and tissue damage, the wavelength of the intense trapping laser must be in the biological transparency window, such as 1064 nm, where residual absorption of tissue components is minimal [10].
In this work, we employed three principal methods to investigate the properties of NV centers in FNDs under NIR light exposure: spin polarization relaxometry (T1-relaxometry) [11, 12], optically detected magnetic resonance (ODMR) spectroscopy [13, 14], and photoluminescence (PL) spectrum analysis [15]. PL spectroscopy involves measuring the emitted light from the NV centers when excited by a laser, providing information on the optical properties and charge states of the NV centers [16]. ODMR combines optical and microwave techniques to detect changes in the NV centers’ fluorescence in response to varying magnetic fields and relative temperature changes, allowing for precise measurement of the NV centers’ electron spin states and local magnetic environment [17]. T1-relaxometry, on the other hand, measures the longitudinal relaxation time (T1), which is the time it takes for the NV centers’ spin states to return to thermal equilibrium after being polarized [18]. This method is valuable for sensing processes and can detect various types of magnetic noise at the spin-transition frequency, such as free radicals [12]. Combined, these techniques provide a comprehensive analysis of the NV centers’ optical, magnetic, and spin properties under varying experimental conditions [4]. We employ a controlled 30 to 60mW NIR laser, with a pulsing sequence inspired by [19], and suggest that the approach can facilitate more precise and reliable intracellular biosensing assessments while keeping any undesirable NIR effect low.
Previous efforts by Aslam et al. [20] have already explored the impact of near-infrared (NIR) laser irradiation on the properties of NV centers, revealing that NIR irradiation can significantly modify NV photoluminescence and enhance spin-lattice relaxation processes, using factors such as NIR laser power (around 30mW) and pulse duration (around 30ns) [21]. The existing literature provides different and sometimes conflicting explanations on the physical mechanisms induced by NIR radiation, which modify both the optical and spin properties of NV centers [22, 15, 9, 21, 23, 14].
In our experiment, we first investigated the effect of the NIR laser on the fluorescence of FNDs in two scenarios: dry FNDs and optically trapped FNDs. We then demonstrated the sensing capabilities of the optically trapped FNDs, which included measurements in different pH environments, controlled temperature variations, and the addition of a paramagnetic species (Gd3+). We propose that changes in spin and photoluminescence properties may be due to three effects: i) heating of the FNDs induced directly by NIR laser irradiation of the FNDs by the NIR laser, ii) a charge state recombination and ionization rate that depends on the NIR laser power and iii) a sensitivity effect, where the NIR laser alters the FNDs’ response to their environment [24]. The sensitivity effect is particularly important for those sensing protocols that rely on NV centers positioned close to the surface of the FNDs [4, 17]. For example, the pH of the surrounding environment plays a significant role in the behavior of NV centers in FNDs. Variations in pH can influence the charge state dynamics of NV centers, which in turn affects their fluorescence properties and overall sensitivity in sensing applications [24, 18, 13, 25].
In this article, we build upon existing hypotheses that the effects of NIR irradiation on NV centers involve both heating, as suggested by previous studies [22], and redistribution between charge states [21, 23]. We aim to provide a more comprehensive understanding of these combined effects, thus offering new insights into optimizing optically trapped NV-based sensing protocols [9]. The paper is organized as follows: Section II describes the methods, including sample preparation, the optical setup, FND trapping, and sensing experiments. Section III presents the results, focusing on photoluminescence spectrum analysis, optically detected magnetic resonance measurements, T1 relaxometry, and a combined statistical analysis. Section IV discusses the implications of the findings, including thermal effects, charge state dynamics, and their interplay, along with alternative considerations. Finally, Section V explores the sensing capabilities of trapped FNDs.
@@ II.1 Sample Preparation
Fluorescent FNDs with a nominal diameter of 120 nm (Sigma-Aldrich, 798088) were used for most of the experiments reported in this article if not stated otherwise. They possess a significant number of nitrogen-vacancy (NV) centers due to high concentration of 3ppm. To prepare the FND suspension, 1 µL of the commercial stock solution of FNDs was diluted in 300-700 µL of deionized water depending on the desired concentration and before sonicating for 10 minutes. Next, the FND suspension was deposited on a petri dish (Mattek, part no P35G-1.5-10-C.), thoroughly cleaned to remove any contaminants that could interfere with the experiments and plasma ionised in order to make the surface hydrophilic, ensuring an even distribution of FNDs on the surface. Plasma treatment was applied for 30 seconds in a simple custom-device with a closed chamber in PMMA and under ambient conditions in which an electrode (Electro Technic Product, BD-20V) could be positioned close to the petri dish surface. The plasma treatment modifies the surface energy of the petri dish by breaking chemical bonds on the surface and introducing polar functional groups, thereby inducing the hydrophilic state [26]. Approximately 3 µL of the FND suspension was then carefully deposited onto the clean petri dish using a micropipette. The suspension was gently spread across the surface before allowing it to dry. The drying process took place at room temperature to minimize any potential alterations to the FND properties potentially caused by excessive heating or cooling. After the solution had completely evaporated, the sample was ready for the optical experiments described in the following. The preparation of FNDs in the liquid environment also starts with the dilution steps and cleaning of the petri dish, but to maintain the FNDs in solution, the petri dish was not plasma treated and the solution was not left for evaporation.
FNDs with a nominal size of 70 nm were used for sensing experiments as they are preferably used in our most recent cell experiments [27]. The FNDs (Sigma-Aldrich 1003626896) were prepared following the same protocol. A suspension of FNDs was created by diluting 1 µL of the commercial stock solution in 1000 µL of deionized water, followed by sonication for 10 minutes to ensure a homogeneous dispersion. Approximately 500 µL of the suspension was deposited in a Petri dish that was thoroughly cleaned.
@@ II.2 Optical Setup
The dried FND sample was placed inside a mini-incubator (pecon XS 2000) in an optical setup based on a customized Thorlabs OTKB/M. Customization included a different objective (Nikon, CFI PLAN APO LAMBDA 100x OIL NA 1.45 WD 0.13) and long-distance condenser lens (Mitutoyo M Plan Apo 10x / 0.26). The setup is constructed around an inverted microscope configuration, equipped with a high-resolution camera (Teledyne Prime 95B), and two lasers for the simultaneous illumination and manipulation of the FNDs. A green laser (Coherent Verdi G5) with a wavelength of 532 nm was utilized for the excitation of the NV centers, enabling the initialization and readout of their spin states. The second laser (Cobolt Rumba, 3W) is operating in the near-infrared (NIR) at a wavelength of 1064 nm and used for FND trapping. This wavelength falls within the biological transparency window, which minimizes absorption and reduces the risk of significant heating effects in biological samples. The two lasers were aligned and their beams combined using a dichroic mirror, ensuring they shared the same optical path before being focused onto the sample through the microscope objective, as detailed in Fig. 2. By trapping an FND and then adjusting the position of the green laser to maximize the signal on the photon counter, we ensured the same focal point for both the green and NIR lasers. Additionally, acousto-optic modulators (AOMs) were used to precisely control the timing and intensity of the laser beams, enabling fine-tuned manipulation of the FNDs.
As illustrated in Fig. 2, several modes of detection were installed: i) camera detection, ii) spectrometer detection and iii) photon count detection. A motorized flip-mirror ensures the ability to rapidly shift between camera detection or either spectrometer or photon count detection. The fluorescence spectrum of the FNDs is recorded and analyzed by a fiber coupled spectrometer (Ocean Optics X00428) while alternatively photon counts are measured with a fiber coupled avalanche photo detector (Excelitas SPCM-AQRH-13-Fc). In addition, to measure the ODMR spectrum of FNDs, an Agilent N5181A signal generator connected to an amplifier (MiniCircuits ZHL-45) is connected to a copper wire antenna placed approximately 1mm above the sample. When the flip-mirror is positioned towards the camera, the latter captures high-resolution images of the FNDs illuminated during the experiments in real time. The combination of different detection methods thus made it possible to continuously monitor the position of the FNDs, determine whether they were isolated or in aggregates, and analyse their photoluminescence properties throughout the experimental process.
Figure 2: Illustration of the optical setup with the most important lasers, optical lenses, filters and equipment.
@@ II.3 FND Trapping
For the FND trapping experiments, solution samples were used to ensure the FNDs remained mobile during the experiment.
The NIR laser was utilized for trapping the FNDs in the focused beam [28, 29, 30, 31]. For $T_{1}$ relaxometry, the temporal sequence of the green and trapping lasers, as illustrated in Fig. 3, is carefully designed to minimize the overlap between the two lasers [19]. This approach reduces potential spin-mixing effects from the NIR laser, while maintaining the stability of the trap. The green laser was pulsed for 35 µs for the excitation and 15 µs for the readout of the NV center spin states. The NIR laser was pulsed for 20 µs for the trapping of the FNDs. By using this sequence, we optimized the trapping stability during $T_{1}$ relaxometry measurement and minimized NIR irradiation potential impact on the measurements.
The sensing experiments utilized the FND trapping protocol described below, and the CW NIR laser power was set to 60 mW to stabilize the FNDs in the optical trap. However, for manipulations involving changes to the surrounding medium, the CW NIR laser power was increased to 100 mW to ensure that the FND remained trapped despite turbulence in the medium. After the modifications, the laser power was reduced back to 60 mW for subsequent measurements.
Figure 3: Laser sequence for T1 measurement. In green the 532 nm laser and in red the 1064 nm laser. $t_{NIR}=20\mu s$, $t_{I}=35\mu s$, $t_{RO}=15\mu s$ and $\tau$ from $7.2\mu s$ to $720\mu s$.
@@ III Results
We begin by presenting the main results of our experiments, starting with the examination of the general fluorescence properties of NV-centers within the FNDs through Photoluminescence (PL) spectroscopy. This foundational assessment provides an overview of the baseline optical characteristics and helps identify any immediate changes induced by the NIR laser irradiation. Following the PL spectrum analysis, we delve into the variations in fluorescence intensity as a function of optically detected magnetic resonance (ODMR) conditions. This step is crucial for understanding the impact of NIR irradiation on the magnetic resonance properties of the NV centers. Finally, we conduct T1-relaxometry to investigate the spin relaxation dynamics, providing a detailed examination of the spin-state populations and their evolution under different NIR laser parameters. This approach from general fluorescence to specific spin dynamics allows for a comprehensive evaluation of the NIR irradiation effects on the NV centers. PL spectrum measurement and ODMR measurement were conducted with CW NIR laser irradiation.

### v2

@@ Abstract
Biosensing based on optically trapped fluorescent nanodiamonds is an intriguing research direction potentially allowing to resolve biochemical processes inside living cells. Towards this goal, we investigate near infrared (NIR) laser irradiation at 1064 nm on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers. The NIR wavelength is a popular choice for optical trapping in biological samples because of its low absorption in bio-samples. By conducting comprehensive experiments, we aim to understand how NIR exposure influences the fluorescence and sensing properties of FNDs and to determine the potential implications for the use of FNDs in various sensing applications. The experimental setup involved exposing FNDs to varying intensities of NIR laser light and analyzing the resultant changes in their optical and physical properties. Key measurements included all-optical fluorescence relaxation, optical spectroscopy, and optically detected magnetic resonance (ODMR) spectra. The findings reveal how increased NIR laser power correlates with alterations in ODMR central frequency but also that charge state dynamics under NIR irradiation of NV centers plays a role. We suggest protocols with NIR and green light that mitigate the effect of NIR, and demonstrate that FND biosensing works well with such a protocol.
@@ I Introduction
Figure 1: Graphical abstract illustrating the sensing modalities explored to investigate effects of NIR laser on NV center for sensing. Temperature, abundance of paramagnetic species, and pH, are altered while optical spectroscopy, ODMR spectra, and fluorescence polarization relaxometry is measured from optically trapped nanodiamonds.
Fluorescent nanodiamonds (FNDs) hosting negatively-charged nitrogen-vacancy (NV-) centers combine bright, photostable fluorescence with quantum-coherent electron-spin states, making them uniquely suited for bio-sensing and bio-imaging at the nanoscale [Schirhagl2014, Liu2021]. Because NV spins can be optically polarised and read out at room temperature, changes in their spin-state populations translate directly into optical signals, enabling quantitative sensing of magnetic fields, temperature and local electro-chemical potentials [Doherty2013, Rondin2014]. The ultimate performance of those sensors, however, is governed by the spin- and charge-state photodynamics of the NV centers themselves; understanding how an experimental control parameter perturbs those dynamics is therefore a prerequisite for accurate, high-sensitivity measurements.
The NV center is an electronic spin-1 system with a ground-state zero-field splitting (ZFS) $D\approx 2.87~\mathrm{GHz}$ that separates the $m_{s}=0$ and $m_{s}=\pm 1$ sub-levels. External magnetic fields lift the degeneracy of the $m_{s}=\pm 1$ levels through the Zeeman effect, while lattice strain, temperature and electric fields shift the ZFS and the optical transition energies via the Stark effect [Maze2008, Dolde2011]. In addition, the relative abundance of the neutral (NV0) and negative (NV-) charge states—and thus the fluorescence spectrum and intensity—depends sensitively on the local Fermi level and on photo-ionisation/recombination pathways. These intertwined dependencies let NV centers act as multi-modal nano-probes, provided that the separate interaction between all external controls (laser beams, microwave fields, etc.) and the FNDs are themselves well understood and do not compromise sensor performance.
A particularly powerful external control is optical trapping. Near-infra-red (NIR) optical tweezers can confine—and manoeuvre—sub-100  nm diamonds inside living cells or complex micro-fluidic environments, offering spatially resolved sensing with a resolution set by the nanoparticle size rather than by the optical diffraction limit [Neukirch2013]. Operating the trap at 1064  nm keeps tissue absorption low and mitigates photo-damage [Svoboda1994], yet the same intense NIR field may also heat the nanodiamond, alter NV charge-state kinetics or modify spin-lattice relaxation. Earlier work on bulk diamond or levitated nanodiamonds indeed reported substantial NIR-induced changes in fluorescence and $T_{\rm relax}$[Aslam2013, Ji2016], but a clear picture for FNDs dispersed in aqueous media—arguably the configuration of greatest practical relevance-is missing.
Here, as illustrated by Fig.1, we address that gap. Using freely diffusing and optically trapped FNDs in water and in biologically relevant buffer solutions, we quantify how a 1064  nm focused trapping beam (30–60  mW average power) affects NV optical and spin properties. We employ tree complementary read-outs:
  1. 1.
    photoluminescence (PL) spectroscopy to monitor charge-state redistribution [Geiselmann2013];
  2. 2.
    optically detected magnetic resonance (ODMR) to track shifts in the resonance frequencies with kHz precision [Gruber1997, Sow2020]; and
  3. 3.
    fluorescence relaxometry to probe magnetic-noise spectra—e.g. from free radicals or $\mathrm{Gd}^{3+}$ ions—at the NV transition frequency [Hall2016].
Together, these observables allow us to disentangle three mechanisms that NIR light can induce:
  1. 1.
    Photothermal heating of the FND, shifting the ZFS via the temperature coefficient $dD/dT\approx-74~\mathrm{kHz\,K^{-1}}$[Acosta2010];
  2. 2.
    Power-dependent ionisation/recombination that re-weights NV- and NV0, thereby changing both ODMR spectral properties and PL brightness [Aslam2013, Ji2016];
  3. 3.
    Modified surface sensitivity, whereby the NIR field—through either heating or altered charge state—amplifies or suppresses the NV response to environmental parameters such as pH or paramagnetic species concentration [Fu2010].
We first benchmark these effects in air-dried FNDs, then in trapped FNDs immersed in solutions of controlled pH and temperature, and finally in the presence of a paramagnetic relaxant ($\mathrm{Gd}^{3+}$). By comparing experiments with and without the NIR trap, we establish quantitative limits on NIR-induced artefacts and delineate operating regimes where accurate, spatially resolved biosensing is possible.
@@ II.1 Sample Preparation
Fluorescent FNDs with a nominal diameter of 120 nm (Sigma-Aldrich, 798088) were used for most of the experiments reported in this article if not stated otherwise. They possess a significant number of nitrogen-vacancy (NV) centers due to a high concentration of 3ppm. To prepare the FND suspension, 1 µL of the commercial stock solution of FNDs was diluted in 300-700 µL of deionized water depending on the desired concentration and before sonicating for 10 minutes. Next, the FND suspension was deposited on a petri dish (Mattek, part no P35G-1.5-10-C.), thoroughly cleaned to remove any contaminants that could interfere with the experiments and plasma ionised in order to make the surface hydrophilic, ensuring an even distribution of FNDs on the surface. Plasma treatment was applied for 30 seconds in a simple custom-device with a closed chamber in PMMA and under ambient conditions. Plasma was generated with a BD-20AV high-frequency generator (230 V, ETP, IL, USA) with an electrode positioned close to the petri dish surface. The plasma treatment modifies the surface energy of the petri dish by breaking chemical bonds on the surface and introducing polar functional groups, thereby inducing the hydrophilic state [Al-Azzam2022]. Approximately 3 µL of the FND suspension was then carefully deposited onto the clean petri dish using a micropipette. The suspension was gently spread across the surface before allowing it to dry. The drying process took place at room temperature to minimize any potential alterations to the FND properties potentially caused by excessive heating or cooling. After the solution had completely evaporated, the sample was ready for the optical experiments described in the following. The preparation of FNDs for experiments in the liquid environment also starts with the dilution steps and cleaning of the petri dish, but to maintain the FNDs in solution, the petri dish was not plasma treated and the solution was not left for evaporation.
FNDs with a nominal size of 70 nm were used for sensing experiments as they are preferably used in our most recent cell experiments [Niora2024]. The FNDs (Sigma-Aldrich 1003626896) were prepared following the same protocol. A suspension of FNDs was created by diluting 1 µL of the commercial stock solution in 1000 µL of deionized water, followed by sonication for 10 minutes to ensure a homogeneous dispersion. Approximately 50 µL of the suspension was deposited in a sterile Petri dish that was freshly unwrapped.
@@ II.2 Optical Setup
The dried FND sample was placed inside a mini-incubator (pecon XS 2000) in an optical setup based on a customized Thorlabs OTKB/M modular optical tweezer system. Customization included a different objective and long-distance condenser lens. Vendors and other details are listed in Table 2 in the Appendix. The setup is constructed around an inverted microscope configuration, equipped with a high-resolution camera that capture wide-field bright-field images of the FND, and two lasers for the simultaneous illumination and manipulation of the FNDs. A green laser with a wavelength of 532 nm was utilized for the excitation of the NV centers, enabling the initialization and readout of their spin states. The second laser is operating in the near-infrared (NIR) at a wavelength of 1064 nm and used for FND trapping. This wavelength falls within the biological transparency window, which minimizes absorption and reduces the risk of significant heating effects in biological samples. The two lasers were aligned and their beams combined using a dichroic mirror, ensuring that they shared the same optical path before being focused onto the sample through the microscope objective, as detailed in Fig. 2a). By trapping an FND and then finetuning the position of the green laser to maximize the signal on the photon counter, we ensured the same focal point for both the green and NIR lasers. Additionally, acousto-optic deflectors were used to precisely control the timing and intensity of the laser beams by using Qudi software, enabling fine-tuned manipulation of the FNDs.
As illustrated in Fig. 2a), several modes of detection were installed: i) camera detection, ii) spectrometer detection and iii) photon count detection. A motorized flip-mirror ensures the ability to rapidly shift between camera detection and either spectrometer or photon count detection. The fluorescence spectrum of the FNDs is recorded and analyzed by a fiber coupled spectrometer while alternatively photon counts are measured with a fiber coupled avalanche photo detector. In addition, to measure the CW ODMR spectrum of FNDs, a signal generator connected to an amplifier is connected to a single-turn copper loop antenna (5 mm) placed approximately 1 mm above the sample. When the flip-mirror is positioned towards the camera, the latter captures high-resolution images of the FNDs illuminated during the experiments in real time. The combination of different detection methods thus made it possible to continuously monitor the position of the FNDs, determine whether they were isolated or in aggregates, and analyse their photoluminescence properties throughout the experimental process.
Figure 2: a) Illustration of the optical setup with the most important lasers, optical lenses, filters and equipment. Details of the components are specified in Table 2 in the appendix. b) Laser sequence for fluorescence relaxometry measurement. In green the 532 nm laser and in red the 1064 nm laser. Pulse-widths are $t_{\rm trap}=20\mu s$, $t_{ini}=35\mu s$, $t_{RO}=15\mu s$ while the variable waiting time $\tau$ allows to record the relaxometry at times spanning from $7.2\mu s$ to $720\mu s$.
@@ II.3 FND Trapping
All trapping experiments were performed in a fluid chamber containing 50 µL of the FND aqueous suspension unless stated otherwise. The NIR laser was utilized for trapping the FNDs in the focused beam [Neuman2004, Gennerich2017, Jones2015]. For relaxometry, the temporal sequence of the green and trapping lasers, as illustrated in Fig. 2b), avoids temporal overlap between the two lasers while maintaining the FND stably trapped [Russell2018a]. Strictly speaking, this protocol does not constitute a conventional spin-relaxometry measurement. With an all-optical protocol, the method inherently involves potential inaccuracies due to photoionization induced by the green laser itself, thus affecting the NV- charge states. To address this limitation, measurement protocols incorporating microwave pulses explicitly designed to disentangle spin dynamics from photoionization effects have been developed. Examples include the microwave $\pi$-pulse method introduced by Ref. [Jarmola2023], as well as alternative sequences described in Ref. [Cardoso2023]. We hence will refer to the measured decay time as $T_{\rm relax}$ throughout the remainder of this paper. The green laser was pulsed with initialization time $t_{ini}$ of 35 µs, and read-out time $t_{RO}$ of 15 µs, for the readout of the NV center fluorescence, while the NIR laser was on to trap the FNDs in sequences with duration $t_{\rm trap}$ of 20 µs.
@@ II.4 Sensing Experiments
For the sensing of paramagnetic species and pH, changes to the surrounding medium was involved. To ensure that the FND remained trapped despite disturbances in the medium, the CW NIR laser power was temporarily increased to 100 mW while liquid was added. Subsequently, the laser power was reduced back to 60 mW for measurements with the new conditions, and the same FND.
@@ III Results
We begin by presenting the main results of our characterisation experiments, starting with the examination of the general fluorescence properties of NV-centers within the FNDs through photoluminescence (PL) spectroscopy. This foundational assessment provides an overview of the baseline optical characteristics and helps identify any immediate changes induced by the NIR laser irradiation. Following the PL spectrum analysis, we present the variations in fluorescence intensity as a function of optically detected magnetic resonance (ODMR) conditions. This step is crucial for understanding the impact of NIR irradiation on the magnetic resonance properties of the NV centers. Finally, we conduct fluorescence relaxometry to investigate the spin relaxation dynamics under different NIR laser parameters. PL spectrum measurement and ODMR measurement were conducted with CW NIR laser irradiation.
The experiments conducted under dry conditions primarily serve as a thermal control benchmark. This dry setup facilitates the understanding of thermal phenomena, such as localized heating effects, by providing a clear reference to compare with conditions involving FNDs suspended in aqueous environments. We acknowledge that the dry condition itself lacks direct practical application in sensing contexts.
For each set of experiments described in the subsections below, we exposed single FNDs to CW NIR laser at power levels of 0 mW, 30 mW, and 60 mW.